OpenStudy (anonymous):

x^2+14x=-48 and x^2+10x=-25

hero (hero):

Hint: x^2 + 14x + 48 = 0 and x^2 + 10x + 25 = 0

OpenStudy (anonymous):

x=-8,-6 for the first one and the last one is x=-5

hero (hero):

Mike if you're going to post an answer, make sure you give complete steps along with explanations

hero (hero):

Answers alone are not welcome here. People can get banned for that sort of thing.

OpenStudy (anonymous):

I did not realize that. Thanks for the suggestion

hero (hero):

Did you read the code of conduct yet?

OpenStudy (anonymous):

Yes I just have read it. Thank you for the suggestion again.

hero (hero):

Hey mike, I'm sure you would also gree that answers alone are not very helpful, right?

hero (hero):

agree*

OpenStudy (anonymous):

I would agree but the asker has no asked for more. Most people here just want to see the answers. Then if they dont agree they might want to see how I got to that answer. Sometimes hints doesn't help well compared to giving the actual answer j/s

hero (hero):

Actually, I'm aware that some (keyword some) users will want "just answers". This is why you should communicate with the student and find out exactly what it is he or she wants. If they say they just want answers, the appropriate response is to say that Open Study's Code of Conduct Policy states that users are not allowed to give answers alone. If the user claims that they are checking their answers, then you should ask them what answers to the problem they have come up with. Open Study does not condone cheating in any way and this is a problem that we've had in the past and are consistently trying to deal with. Open Study should be a place where people come to study or get help, not get just answers. If they wanted Open Study to be about getting answers, they would have named it "Open Answer".

OpenStudy (anonymous):

Ok once again thanks for the suggestion no need to be trolling.

hero (hero):

I'm just here to help you mike.

hero (hero):

If I wanted to troll, I would not be here.

OpenStudy (anonymous):

Like I said again thanks for the suggestion for the 5th time

hero (hero):

I was only responding to your previous comments where it seemed you needed further clarification on the matter. I'm sorry if it seems like I'm being redundant.